OpenStudy (anonymous):

When the smaller of two consecutive integers is added to four times the larger, the result is 29. Find the integers

OpenStudy (anonymous):

Let small number = x Let large number = y x+4y=29 y-1=x or x+1=y

OpenStudy (anonymous):

It is a system, solve for x and y

OpenStudy (anonymous):

do you know how to?

Directrix (directrix):

Let x be an integer (smaller) and x + 1 the consecutive integer (larger) smaller of two consecutive integers is added to four times the larger So, x is added to 4*(x + 1) to give x + 4*(x+1) The sum is 29. Putting that together, x + 4*(x+1) = 29 x + 4x + 4 = 29 5x + 4 = 29 -------->@kaela_stanley --> Solve this for x. To that, add 1 to get the larger integer.